OpenStudy (anonymous):

3x^3+2x^2 +X+5 the derivative evalauate y' when x=3 need a little help bit rusty

OpenStudy (anonymous):

in order to the first you lower the power by one like this 3x^2+2x^1+5 that is der right?

OpenStudy (unklerhaukus):

you also have to divide by the original index

OpenStudy (anonymous):

you divide by 3 because of 3x^3

OpenStudy (anonymous):

x^n=nx^n-1

OpenStudy (anonymous):

so if I divide by 3, than that would make this x^2 instead of 3x^2 I am on the right track?

OpenStudy (unklerhaukus):

opps ive got mixed up

OpenStudy (anonymous):

y' 3x^2 is X^2 right?

OpenStudy (unklerhaukus):

i should have said times by the index \[y(x)=ax^n\qquad\Rightarrow\qquad y'(x)=anx^{n-1}\]

OpenStudy (unklerhaukus):

so sorry if i confused you

OpenStudy (anonymous):

so times by index makes 3x^3 Y' 6x^2 right?

OpenStudy (unklerhaukus):

not quite \[y(x)=3x^3\qquad y'(x)=3\times3x^2\]

OpenStudy (anonymous):

here is what I know y' when doing this you lower the power by one and something??

OpenStudy (unklerhaukus):

when taking the derivative multiply by the power and then lower the power

OpenStudy (anonymous):

so you are telling that power before lowering divide to y'

OpenStudy (anonymous):

like this 4x^3 so that would be 'y 4 right?

OpenStudy (anonymous):

4x

OpenStudy (unklerhaukus):

i ment times when i said divide originally , dont divide when taking the derivative

OpenStudy (anonymous):

can we start over plz..

OpenStudy (unklerhaukus):

yes ,

OpenStudy (anonymous):

y' 4x^2 would y'4x

OpenStudy (unklerhaukus):

\[y(x)=ax^n\qquad\Rightarrow\qquad y'(x)=n\times ax^{n-1}\] \[y(x)=4x^2\qquad\Rightarrow\qquad y'(x)=2\times 4x^{2-1}\]

OpenStudy (anonymous):

y'(x) would be 8x

OpenStudy (unklerhaukus):

thats right now

OpenStudy (anonymous):

okay so the power is what you are times before lowering the power?

OpenStudy (unklerhaukus):

that is right

OpenStudy (anonymous):

so for another example 10X^5 y'(x) 50x^4

OpenStudy (unklerhaukus):

you got it \(\checkmark\)

OpenStudy (anonymous):

so back to the problem now

OpenStudy (unklerhaukus):

\[y(x)=3x^3+2x^2 +x+5\]\[y'(x)=\]

OpenStudy (anonymous):

y'(x) 6x+4X +5

OpenStudy (unklerhaukus):

not quite, note that \[x=x^1\]\[5=5x^0\]

OpenStudy (anonymous):

y'(x) 9x+4+5

OpenStudy (unklerhaukus):

try again

OpenStudy (anonymous):

y'(x) 6x^2 +4x+5 +1

OpenStudy (unklerhaukus):

ok lets do this one term at a time \[(3x^3)^\prime=\]

OpenStudy (anonymous):

y'(x) 9x^2

OpenStudy (unklerhaukus):

right, now \[(2x^2)^\prime=\]

OpenStudy (anonymous):

y'(x) 9x^2 +2x

OpenStudy (anonymous):

y'(x) 9x^2 4+5+1

OpenStudy (unklerhaukus):

nope

OpenStudy (anonymous):

y'(x) 9x^2 +4X+5

OpenStudy (unklerhaukus):

you have the first two terms right,

OpenStudy (unklerhaukus):

\[f(x)=x=x^1\] \[f^\prime(x)=1\times x^{1-1}=x^0=1\]

OpenStudy (anonymous):

so y'(x) 9x^2+4x+1+5

OpenStudy (unklerhaukus):

the derivative of the last term \[g(x)=5=5x^0\] \[g^\prime(x)=0\times 5x^{0-1}=\]

OpenStudy (anonymous):

is 0

OpenStudy (unklerhaukus):

yeah the derivative of a constant is always zero

OpenStudy (anonymous):

so the end number unless it has a power drop off?

OpenStudy (unklerhaukus):

thats right

OpenStudy (unklerhaukus):

\[f(x)=x^2+5\qquad f'(x)=2x\]\[g(x)=x^2+9\qquad g'(x)=2x\]\[h(x)=x^2−71\qquad f'(x)=2x\]

OpenStudy (unklerhaukus):

that last one is ment to be h'

OpenStudy (anonymous):

cool love it so far because in math you can't drop anything, in calc 1 you can this is way cool..

OpenStudy (anonymous):

do you have time to do one more?

OpenStudy (unklerhaukus):

yeah can you get the answer to you original question now?

OpenStudy (unklerhaukus):

\[y(x)=3x^3+2x^2 +x+5\] \[y~'(x)=\]

OpenStudy (anonymous):

y'(x) 9x^2+4x+1

OpenStudy (unklerhaukus):

\[\Large\color\red\checkmark\]

OpenStudy (anonymous):

okay any other tips here?

OpenStudy (unklerhaukus):

\[y'(x)=\frac{\text dy(x)}{\text dx}\]

OpenStudy (anonymous):

explain a little??

OpenStudy (unklerhaukus):

its just a different notation, they men the same thing

OpenStudy (unklerhaukus):

the second has a bit more information because it explicitly say that we are to take the derivative of the function with respect to x, ie we look at the power of the x terms,

OpenStudy (anonymous):

dx(x)^1?dx

OpenStudy (anonymous):

so the power on the is 1 right?

OpenStudy (anonymous):

just to let know, I am in college not high school

OpenStudy (unklerhaukus):

\[y(x)=x=x^1\] \[\frac{\text dy(x)}{\text dx}=y'(x)=1\times x^{1-1}=x^0=1\]

OpenStudy (unklerhaukus):

im not sure what you mean by this dx(x)^1?dx

OpenStudy (anonymous):

I meant to use / not?

OpenStudy (unklerhaukus):

dx(x)^1/dx \[=\large \frac{\text d(x)^1}{\text dx}\]?

OpenStudy (anonymous):

yes that is what I meant to type

OpenStudy (unklerhaukus):

\[\frac{\text d(x)}{\text dx}=\frac{\text d(x^1)}{\text dx}=1\times x^0=1\]

OpenStudy (unklerhaukus):

im not sure if i am answer your question?

OpenStudy (anonymous):

okay now for the last question here how do this with y"(x) sin

OpenStudy (anonymous):

yes you have

OpenStudy (unklerhaukus):

should we evaluate y' in the first question when x=3 before that?

OpenStudy (anonymous):

y'(x) 9x^2+4x+1 so plug in 3 so we have y'(x) 9x^2 (3) +4x(3) +1 which is y'(x) 27x^2+ 12X +1

OpenStudy (unklerhaukus):

\[y(x)=3x^3+2x^2+x+5\] \[y'(x)=9x^2+4x+1\] \[y'(3)=9(3)^2+4(3)+1=9\times3\times3+4\times3+1\]

OpenStudy (anonymous):

y'(3) 81x^2+12+1

OpenStudy (anonymous):

12x

OpenStudy (unklerhaukus):

when you evaluate a function at a particular x value you substitute this value for x in the equation, there should be no x in left when evaluated

OpenStudy (anonymous):

sorry i mess me up by not putting the like this (3)^2

OpenStudy (anonymous):

so the answer is F'(3) 81x^2 + 12X +1